# Dirac Fermions and Topological Phases in Magnetic Topological Insulator Films

### v2

@@ V.2.2 Asymmetric magnetic structure
Pictures and discussions above complete the case study for the Chern insulator phase here. Notice that in the typical cases given above, the Zeeman field directs along $z$-positive axis, and it is always the $\chi=-$ band that has $-e^{2}/h$ Hall conductivity while the $\chi=+$ band is trivial with zero Hall contribution, i.e., it is a $1+0$ combination with the sign of Hall conductivity determined by the polarization direction of the Zeeman field, as we shall illustrate further below.
@@
$\begin{cases}C_{-}=1,\leavevmode\nobreak\ C_{+}=0,&\text{ for }I_{S}(0)>0\\ C_{+}=-1,\leavevmode\nobreak\ C_{-}=0,&\text{ for }I_{S}(0)<0\end{cases},$ (102)
@@
we find that for the asymmetric Chern insulator case $-\chi I_{S}(0)>|I_{A}(0)|\geq 0$, and it always holds that
@@ V.2.3 Mirror layer Chern number
$H_{\chi}=P^{M_{z}}_{\chi}H,\leavevmode\nobreak\ P^{M_{z}}_{\chi}=\frac{1+i\chi M_{z}}{2},$
with $M_{z}$ the represented mirror operator, and its anti-diagonal elements are recognized to be $U_{z}$, which relates quantity at ±$\pm l_{z}$ (see section V.1 for more about mirror symmetry).
@@
where in the usual case, the expectation value of velocity operator is $\bar{v}_{i}^{mn}(l)=\braket{m(l)}{\partial_{k_{i}}H}{n(l)}$ with only diagonal elements, which, however, fails for the mirror projected Hamiltonian. The key observation lies in the fact that by projection $\partial_{k_{i}}H_{\chi}$ contains not only diagonal elements but off-diagonal part, which induces additional non-local transition contribution from exactly mirror symmetrized layers. Work the effect out and one obtains the mirror layer Hall conductivity
@@
Figure 16: Mirror layer Hall conductivity for topological insulator film with symmetric $z$-Zeeman field immersed at top $2$ and bottom $2$ layers upon total $19$ layers with strength $V_{z}=0.1$\mathrm{e}\mathrm{V}$$, and chemical potential is chosen to be $\mu=2.5$\mathrm{m}\mathrm{e}\mathrm{V}$$, where (a) for $\chi=-$, (b) for $\chi=+$ and (c) for the total result by adding two mirror parts together, respectively. Both layer-resolved and layer-added Hall conductivity are presented. To respect the mirror symmetry we put the TI film at origin and the layer index becomes $l_{z}=-\frac{L_{z}-1}{2},-\frac{L_{z}-3}{2},\cdots,\frac{L_{z}-1}{2}$ with total layer number $L_{z}=19$.
@@
The layer dependent Hall conductivity serves us a new insight to understand the phenomenon. Treating the system as a whole, its layer-resolved Hall conductivity, as presented in Fig. 16(c), becomes non-zero mainly near the top and bottom surfaces where time-reversal symmetry is broken explicitly under the Zeeman field. And the cumulated Hall conductivity gains approximately half quantum Hall conductivity near two surfaces. On the other hand, as shown in Fig. 16(a), (b), when we split the system by mirror symmetry, the layer-resolved mirror Hall conductivity shows similar top and bottom distribution as the whole system, but with only half the amplitude by mirror splitting, while the Hall conductivity distribution around mirror plane shows opposite-sign peaks inherited from the time-reversal unbroken bulk property like that in the half quantum mirror Hall effect. Once the Hall conductivity contribution is added layer by layer, we immediately see the tri-section configuration: for the non-trivial $C_{-}=1$ part, there exist two Hall-plateaus separating the surface and bulk, then following the top-middle-bottom section cut, we see a contribution rather close to $(-1/4)$$(-1/2)$$(-1/4)$ from each section; and for the trivial $C_{+}=0$ part, the section separation is not that apparent, and we only roughly write $(-c/4)$$(c/2)$$(-c/4)$ with $c$ approximately one to represent the observed distribution.
@@ V.3 Axion insulator: An antisymmetric magnetic structure
Figure 18: Axion insulator: total layer number $L_{z}=19$ with top-$2$-layer Zeeman field $V_{z}^{t}=0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ and bottom-$2$-layer field $V_{z}^{b}=-0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.
@@ V.4 MnBi2Te4 film: Even and odd number of magnetic layers
Figure 20: Left (right) pictures are for $L_{z}=19\leavevmode\nobreak\ (18)$ anti-ferromagnetic TI film as an odd (even) one. The Zeeman strength is chosen to be $|V_{z}|=0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$. (a) (b) Calculated $I_{S/A}(\bm{k})$ for the effective model. (c) (d) Calculated Hall conductance from magnetic TI film Hamiltonian.
@@ V.4.1 Odd layer: Chern insulator
$\begin{cases}I_{S}>0\\ I_{A}=0\end{cases},\leavevmode\nobreak\ L_{z}\mod 2=1,$ (110)
@@ V.4.2 Even layer: Axion insulator
$\begin{cases}I_{S}=0\\ I_{A}>0\end{cases},\leavevmode\nobreak\ L_{z}\mod 2=0,$ (111)
@@ V.5 Half-quantized anomalous Hall effect: A semi-magnetic film
Figure 22: Half-quantized anomalous Hall metal: total layer number $L_{z}=19$ with top-$2$-layer Zeeman field $V_{z}^{t}=0.1\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductance from TI film model and effective four-band Hamiltonian. (d) Hall conductance for $\chi=\pm$.
@@
The gapless Dirac fermion, identified as the band with gapless surface states contributes a half-quantized Hall conductance. From Fig. 22 (d), the $\chi=+$ band is trivial with zero-Hall plateau inside the Zeeman gap, i.e., the Zeeman gapped band is trivial, while the $\chi=-$ band contains a relatively large Hall plateau quantized to $-e^{2}/2h$, which is bounded by the TI bulk gap and corresponds to the Hall conductance contributed from the high-energy part of the gapless Dirac band[31]. To explain this behavior, it is important to note that we now have $I\equiv I_{S}=I_{A}>0$ around the Dirac point revealed in Fig. 22 (b) (valid in the parity invariant regime bounded by $k_{c}$), and the effective masses become, according to Eq. (81),
@@
$\begin{cases}\sigma_{H}^{\chi}=\dfrac{\chi}{2}\dfrac{e^{2}}{h}\\ \sigma_{H}^{\bar{\chi}}=0\end{cases},\leavevmode\nobreak\ |\mu|<2|I_{A}(0)|,$ (117)
@@
counting on the fact that $m(\bm{k})\leq 0$. This effective Hamiltonian offers with half-quantized Hall conductance $-\mathrm{sgn}(V)e^{2}/2h$, which does not depend on whether the magnetism is put at the top or bottom of TI film, but only on its polarization direction. Under an external magnetic field, such a single gapless Dirac fermion will step into the quantum Hall regime[90, 52] and exhibits quantized Hall conductance whenever an integer number of Landau levels become fully filled[139]. Especially, the ‘anomaly’ contribution will manifest itself to compensate the half quantization contributed from the lowest Landau level, so as to keep the integer value of Chern invariant for this gapped Landau level system.
@@ VI.1 Metallic quantized anomalous Hall effect: A film with a magnetic sandwich structure
Figure 26: The evolution of the effective mass $\tilde{m}_{n,\chi}(k_{x},k_{y}=0)$ ($n=1,2$). (a)$\sim$(g) and (h)$\sim$(n) show lowest-two effective masses varying with changing Zeeman field strength $\alpha$ belonging to set $[0,0.3,0.5,0.7,0.74,0.8,0.9]$ for $\chi=+$ and $\chi=-$, respectively. (a) (g) (h) (n) have already been shown as Fig. 3 in the main text but with a finer structure here. Adapted from [32].
@@
$V_{z}(l_{z})=\begin{cases}\alpha t_{\perp},\leavevmode\nobreak\ &l_{z}=\pm 1/2,\cdots,\pm(m_{z}-1)/2\\ 0,&\text{otherwise}\end{cases},$ (125)
@@
Attributed to the mass exchange mechanism over the effective mass picture presented in Section IV, such a topological phase transition with the increasing of $\alpha$ as Zeeman strength in the middle can be explained. Absorbing the $\alpha$-dependent Zeeman term into the one-dimensional Hamiltonian separated from the TI film leads to an $\alpha$-dependent 1-D Hamiltonian $H_{1d}(\alpha)$, with $H_{1d}(\alpha=0)$ coming back to the 1-D Hamiltonian extracted from TI film and solved exactly before (see section II.2). Projecting $H_{1d}(\alpha)$ over solutions of $H_{1d}(0)$ leads to $\left(\bigoplus_{n=1}^{L_{z}}m_{n}\tau_{z}+\alpha\mathbf{I}_{S}(\mathbf{k})\tau_{0}\right)\sigma_{z}$, and further diagonalizing this provides a bijection which maps the projected Hamiltonian form into the mass term $\bigoplus_{n=1,\chi=\pm}^{L_{z}}\tilde{m}_{n,\chi}(\mathbf{k},\alpha)\sigma_{z}$ (see section IV.1). Notice that both $\sigma_{z}$ and $\tau_{z}$ here are good quantum numbers, as spin and mirror indices ($\chi=\pm$), respectively. Confining to the subspace with $\sigma_{z}=+$, we could then track the evolution and interaction of the mass terms $\tilde{m}_{n,\chi}$ between $n=1$ and $n=2$ blocks with increasing $\alpha$ for given $\chi$. As shown in Fig. 26, $\tilde{m}_{n,+}(n=1,2)$ maintain their shapes increasing $\alpha$, while $\tilde{m}_{n,-}(n=1,2)$ have effectively exchanged their high-energy parts through the low-energy mass exchange, which leads to the high-energy mass sign change of the gapless Dirac cone, and alters its Hall conductivity from $e^{2}/2h$ to $-e^{2}/2h$, when Fermi surface lies inside the parity invariant regime. Then combined with the unaltered $-e^{2}/2h$ from $\tilde{m}_{1,+}$, totally a topological phase transition happens, driving the system from zero Hall conductivity to quantized Hall conductivity, with Hall contribution coming from two metallic bands, which makes the system a metallic topological phase. We can identify
@@ VI.1.1 A qualitatively model with $n=1,2$
$\tilde{m}_{\chi,\zeta}=(E_{\chi}(\bm{k})+\zeta\Lambda_{\chi}(\bm{k}))\sigma_{z},\leavevmode\nobreak\ \chi,\zeta=\pm,$ (133)
@@
$\begin{cases}m_{1}(0)=0,\leavevmode\nobreak\ m_{2}(0)>0\\ -m_{1}(M)\approx m_{2}(M)\gg\alpha|I_{S}(M)|>0\\ I_{S}^{11}(0)=I_{S}^{12}(0)=0,\leavevmode\nobreak\ I_{S}^{22}(0)>0\\ I_{A}=0\end{cases},$
@@
with total Hall conductivity unit upon $e^{2}/h$. This unit is fundamentally different the $C=1$ as Chern insulator case, since here $1=1/2+1/2$, with non-vanishing contribution coming from two metallic bands describing gapless Dirac fermions. It is recognized that the phase transition happens only within $\chi=-$ sub-blocks, where $\zeta=\pm$ Dirac fermions exchange their low-energy mass when crossing the qualitative phase transition point $I_{S}^{22}(0)=m_{2}(0)$, and by treating approximately $I_{S}^{22}\approx\alpha t_{\perp},\leavevmode\nobreak\ m_{2}(0)\approx m_{0}$, we see the qualitative critical point as
@@ VI.1.2 Lower threshold by decreasing the mass in the middle
Figure 28: The evolution of the effective mass $\tilde{m}_{n,\chi}(k_{x},k_{y}=0)$. (a)$\sim$(g) and (h)$\sim$(n) emphasize lowest-two ($n=1,2$) effective masses in red and blue colors, varying with changing Zeeman field strength $\alpha$ belonging to set $[0,0.1,0.25,0.4,0.435,0.5,0.6]$ for $\chi=+$ and $\chi=-$, respectively. Still, we take total layer number $L_{z}=22$ and middle Zeeman layer number $m_{z}=6$, while the difference with Fig. 26 is that here the middle-layer mass is reduced to $\tilde{m}_{0}=0.08\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$.
It was pointed out[141, 142] that magnetic doping can reduce and even drive the bulk band gap $m_{0}$ of TI into a trivial one, and this effect plays a positive role in realizing the metallic QAHE indeed. To illustrate this, consider a simplified scenario where the bulk mass of TI, initially $m_{0}=0.28\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$, is reduced to $\tilde{m}_{0}=0.08\leavevmode\nobreak\ $\mathrm{e}\mathrm{V}$$ in the magnetically doped region. Then by comparing the detailed effective mass evolution in Fig. 28 with the original case in Fig. 26, we observe that the critical point $\alpha_{c}$ decreases to approximately $\alpha_{c}\approx 0.435$. Such a reduction is beneficial for the experimental realization of the metallic QAHE. Moreover, since the decrease in the critical $\alpha$ is positively correlated with the reduction of doped middle layer mass, while and this mass reduction itself is also positively correlated with the increase in concentration of magnetic doping, it is expected that the metallic QAHE can be achieved with a significantly lower threshold of magnetic doping concentration in practice. Note that if the middle layers are driven to a trivial state with a negative bulk mass $\tilde{m}_{0}<0$, and are simultaneously considered nonconductive, the system effectively splits into two semi-magnetic TI films, a trivial metallic QAHE comprising two non-communicative half QAHEs with the same half-quantized Hall conductance is obtained. It is important to recognize that the above calculation assumes an oversimplified relationship between doping concentration and the reduced mass. A more accurate determination of the modified critical point requires a realistic model and a self-consistent calculation.

### v3

@@ V.2.2 Asymmetric magnetic structure
Pictures and discussions above complete the case study for the Chern insulator phase here. Notice that in the typical cases given above, the Zeeman field directs along $z$-positive axis, and it is always the $\chi=$- band that has $e^{2}/h$ Hall conductivity while the $\chi=+$ band is trivial with zero Hall contribution, i.e., it is a $1+0$ combination with the sign of Hall conductivity determined by the polarization direction of the Zeeman field, as we shall illustrate further below.
@@
[…]
@@
we find that for the asymmetric Chern insulator case $\chi I_{S}(0)>|I_{A}(0)|\geq 0$, and it always holds that
@@ V.2.3 Mirror layer Chern number
$H_{\chi}=P^{M_{z}}_{\chi}H,\penalty 10000\ P^{M_{z}}_{\chi}=\frac{1+i\chi M_{z}}{2},$
with $M_{z}$ the represented mirror operator, and its anti-diagonal elements are recognized to be $U_{z}$, which relates quantity at $\pm l_{z}$ (see section V.1 for more about mirror symmetry).
@@
where in the usual case, the expectation value of velocity operator is $\bar{v}_{i}^{mn}(l)=\braket{m(l)|\partial_{k_{i}}H|n(l)}$ with only diagonal elements, which, however, fails for the mirror projected Hamiltonian. The key observation lies in the fact that by projection $\partial_{k_{i}}H_{\chi}$ contains not only diagonal elements but off-diagonal part, which induces additional non-local transition contribution from exactly mirror symmetrized layers. Work the effect out and one obtains the mirror layer Hall conductivity
@@
Figure 16: Mirror layer Hall conductivity for topological insulator film with symmetric $z$-Zeeman field immersed at top $2$ and bottom 22 layers upon total 1919 layers with strength $V_{z}=0.1$\mathrm{e}\mathrm{V}$\$, and chemical potential is chosen to be $\mu=2.5$\mathrm{m}\mathrm{e}\mathrm{V}$\$, where (a) for $\chi=$-, (b) for $\chi=+$ and (c) for the total result by adding two mirror parts together, respectively. Both layer-resolved and layer-added Hall conductivity are presented. To respect the mirror symmetry we put the TI film at origin and the layer index becomes $l_{z}=-\frac{L_{z}-1}{2},-\frac{L_{z}-3}{2},\cdots,\frac{L_{z}-1}{2}$ with total layer number $L_{z}=19$.
@@
The layer dependent Hall conductivity serves us a new insight to understand the phenomenon. Treating the system as a whole, its layer-resolved Hall conductivity, as presented in Fig. 16(c), becomes non-zero mainly near the top and bottom surfaces where time-reversal symmetry is broken explicitly under the Zeeman field. And the cumulated Hall conductivity gains approximately half quantum Hall conductivity near two surfaces. On the other hand, as shown in Fig. 16(a), (b), when we split the system by mirror symmetry, the layer-resolved mirror Hall conductivity shows similar top and bottom distribution as the whole system, but with only half the amplitude by mirror splitting, while the Hall conductivity distribution around mirror plane shows opposite-sign peaks inherited from the time-reversal unbroken bulk property like that in the half quantum mirror Hall effect. Once the Hall conductivity contribution is added layer by layer, we immediately see the tri-section configuration: for the non-trivial $C_{-}=1$ part, there exist two Hall-plateaus separating the surface and bulk, then following the top-middle-bottom section cut, we see a contribution rather close to (1/4)(-1/4)(1/2)(-1/2)(1/4)(-1/4) from each section; and for the trivial $C_{+}=0$ part, the section separation is not that apparent, and we only roughly write (c/4)(-c/4)(c/2)(c/2)(c/4)(-c/4) with $c$ approximately one to represent the observed distribution.
@@ V.3 Axion insulator: An antisymmetric magnetic structure
Figure 18: Axion insulator: total layer number $L_{z}=19$ with top-22-layer Zeeman field $V_{z}^{t}=0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$ and bottom-22-layer field $V_{z}^{b}=-0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductivity from TI film model and effective four-band Hamiltonian. (d) Hall conductivity for $\chi=\pm$.
@@ V.4 MnBi2Te4 film: Even and odd number of magnetic layers
Figure 20: Left (right) pictures are for $L_{z}=19\penalty 10000\ (18)$ anti-ferromagnetic TI film as an odd (even) one. The Zeeman strength is chosen to be $|V_{z}|=0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$. (a) (b) Calculated $I_{S/A}(\bm{k})$ for the effective model. (c) (d) Calculated Hall conductance from magnetic TI film Hamiltonian.
@@ V.4.1 Odd layer: Chern insulator
$\begin{cases}I_{S}>0\\ I_{A}=0\end{cases},\penalty 10000\ L_{z}\mod 2=1,$ (110)
@@ V.4.2 Even layer: Axion insulator
$\begin{cases}I_{S}=0\\ I_{A}>0\end{cases},\penalty 10000\ L_{z}\mod 2=0,$ (111)
@@ V.5 Half-quantized anomalous Hall effect: A semi-magnetic film
Figure 22: Half-quantized anomalous Hall metal: total layer number $L_{z}=19$ with top-22-layer Zeeman field $V_{z}^{t}=0.1\penalty 10000\ $\mathrm{e}\mathrm{V}$\$. (a) Comparison of band structure from the lowest four bands of TI film model and effective four-band Hamiltonian. (b) Calculated $I_{S}(\bm{k})$ and $I_{A}(\bm{k})$. (c) Calculated Hall conductance from TI film model and effective four-band Hamiltonian. (d) Hall conductance for $\chi=\pm$.
@@
The gapless Dirac fermion, identified as the band with gapless surface states contributes a half-quantized Hall conductance. From Fig. 22 (d), the $\chi=+$ band is trivial with zero-Hall plateau inside the Zeeman gap, i.e., the Zeeman gapped band is trivial, while the $\chi=$- band contains a relatively large Hall plateau quantized to $e^{2}/2h$, which is bounded by the TI bulk gap and corresponds to the Hall conductance contributed from the high-energy part of the gapless Dirac band[31]. To explain this behavior, it is important to note that we now have $I\equiv I_{S}=I_{A}>0$ around the Dirac point revealed in Fig. 22 (b) (valid in the parity invariant regime bounded by $k_{c}$), and the effective masses become, according to Eq. (81),
@@
$\begin{cases}\sigma_{H}^{\chi}=\dfrac{\chi}{2}\dfrac{e^{2}}{h}\\ \sigma_{H}^{\bar{\chi}}=0\end{cases},\penalty 10000\ |\mu|<2|I_{A}(0)|,$ (117)
@@
counting on the fact that $m(\bm{k})\leq 0$. This effective Hamiltonian offers with half-quantized Hall conductance $\mathrm{sgn}(V)e^{2}/2h$, which does not depend on whether the magnetism is put at the top or bottom of TI film, but only on its polarization direction. Under an external magnetic field, such a single gapless Dirac fermion will step into the quantum Hall regime[90, 52] and exhibits quantized Hall conductance whenever an integer number of Landau levels become fully filled[139]. Especially, the ‘anomaly’ contribution will manifest itself to compensate the half quantization contributed from the lowest Landau level, so as to keep the integer value of Chern invariant for this gapped Landau level system.
@@ VI.1 Metallic quantized anomalous Hall effect: A film with a magnetic sandwich structure
[…]
@@
$V_{z}(l_{z})=\begin{cases}\alpha t_{\perp},\penalty 10000\ &l_{z}=\pm 1/2,\cdots,\pm(m_{z}-1)/2\\ 0,&\text{otherwise}\end{cases},$ (125)
@@
Attributed to the mass exchange mechanism over the effective mass picture presented in Section IV, such a topological phase transition with the increasing of $\alpha$ as Zeeman strength in the middle can be explained. Absorbing the $\alpha$-dependent Zeeman term into the one-dimensional Hamiltonian separated from the TI film leads to an $\alpha$-dependent 1-D Hamiltonian $H_{1d}(\alpha)$, with $H_{1d}(\alpha=0)$ coming back to the 1-D Hamiltonian extracted from TI film and solved exactly before (see section II.2). Projecting $H_{1d}(\alpha)$ over solutions of $H_{1d}(0)$ leads to $\left(\bigoplus_{n=1}^{L_{z}}m_{n}\tau_{z}+\alpha\mathbf{I}_{S}(\mathbf{k})\tau_{0}\right)\sigma_{z}$, and further diagonalizing this provides a bijection which maps the projected Hamiltonian form into the mass term $\bigoplus_{n=1,\chi=\pm}^{L_{z}}\tilde{m}_{n,\chi}(\mathbf{k},\alpha)\sigma_{z}$ (see section IV.1). Notice that both $\sigma_{z}$ and $\tau_{z}$ here are good quantum numbers, as spin and mirror indices ($\chi=\pm$), respectively. Confining to the subspace with $\sigma_{z}=+$, we could then track the evolution and interaction of the mass terms $\tilde{m}_{n,\chi}$ between $n=1$ and $n=2$ blocks with increasing $\alpha$ for given $\chi$. As shown in Fig. 26, $\tilde{m}_{n,+}(n=1,2)$ maintain their shapes increasing $\alpha$, while $\tilde{m}_{n,-}(n=1,2)$ have effectively exchanged their high-energy parts through the low-energy mass exchange, which leads to the high-energy mass sign change of the gapless Dirac cone, and alters its Hall conductivity from $e^{2}/2h$ to $e^{2}/2h$, when Fermi surface lies inside the parity invariant regime. Then combined with the unaltered $e^{2}/2h$ from $\tilde{m}_{1,+}$, totally a topological phase transition happens, driving the system from zero Hall conductivity to quantized Hall conductivity, with Hall contribution coming from two metallic bands, which makes the system a metallic topological phase. We can identify
@@ VI.1.1 A qualitatively model with $n=1,2$
$\tilde{m}_{\chi,\zeta}=(E_{\chi}(\bm{k})+\zeta\Lambda_{\chi}(\bm{k}))\sigma_{z},\penalty 10000\ \chi,\zeta=\pm,$ (133)
@@
$\begin{cases}m_{1}(0)=0,\penalty 10000\ m_{2}(0)>0\\ -m_{1}(M)\approx m_{2}(M)\gg\alpha|I_{S}(M)|>0\\ I_{S}^{11}(0)=I_{S}^{12}(0)=0,\penalty 10000\ I_{S}^{22}(0)>0\\ I_{A}=0\end{cases},$
@@
with total Hall conductivity unit upon $e^{2}/h$. This unit is fundamentally different the $C=1$ as Chern insulator case, since here $1=1/2+1/2$, with non-vanishing contribution coming from two metallic bands describing gapless Dirac fermions. It is recognized that the phase transition happens only within $\chi=$- sub-blocks, where $\zeta=\pm$ Dirac fermions exchange their low-energy mass when crossing the qualitative phase transition point $I_{S}^{22}(0)=m_{2}(0)$, and by treating approximately $I_{S}^{22}\approx\alpha t_{\perp},\penalty 10000\ m_{2}(0)\approx m_{0}$, we see the qualitative critical point as
@@ VI.1.2 Lower threshold by decreasing the mass in the middle
[…]
It was pointed out[141, 142] that magnetic doping can reduce and even drive the bulk band gap $m_{0}$ of TI into a trivial one, and this effect plays a positive role in realizing the metallic QAHE indeed. To illustrate this, consider a simplified scenario where the bulk mass of TI, initially $m_{0}=0.28\penalty 10000\ $\mathrm{e}\mathrm{V}$\$, is reduced to $\tilde{m}_{0}=0.08\penalty 10000\ $\mathrm{e}\mathrm{V}$\$ in the magnetically doped region. Then by comparing the detailed effective mass evolution in Fig. 28 with the original case in Fig. 26, we observe that the critical point $\alpha_{c}$ decreases to approximately $\alpha_{c}\approx 0.435$. Such a reduction is beneficial for the experimental realization of the metallic QAHE. Moreover, since the decrease in the critical $\alpha$ is positively correlated with the reduction of doped middle layer mass, while and this mass reduction itself is also positively correlated with the increase in concentration of magnetic doping, it is expected that the metallic QAHE can be achieved with a significantly lower threshold of magnetic doping concentration in practice. Note that if the middle layers are driven to a trivial state with a negative bulk mass $\tilde{m}_{0}<0$, and are simultaneously considered nonconductive, the system effectively splits into two semi-magnetic TI films, a trivial metallic QAHE comprising two non-communicative half QAHEs with the same half-quantized Hall conductance is obtained. It is important to recognize that the above calculation assumes an oversimplified relationship between doping concentration and the reduced mass. A more accurate determination of the modified critical point requires a realistic model and a self-consistent calculation.